OpenStudy (anonymous):

What is the slope of the line 2x + 1/3 y = –8?

OpenStudy (kanwal32):



OpenStudy (anonymous):

is it \[\frac{ 2x+1 }{ 3y } = -8 ????????\]

OpenStudy (anonymous):

A. –6 B. –24 C. 2 D. -1/3 these are my options.

OpenStudy (anonymous):

i am asking about the format of the ques is it \[\frac{ 2x+1 }{ 3y } = -8 ?????\] OR \[2x +\frac{ 1 }{ 3y } = -8\]????

OpenStudy (anonymous):

the y is right after 1/3